### v1

@@ 4 Discussion
In this study, we analyzed EEG data from a cohort of 87 participants exposed to a stressful work-related scenario presented through a virtual reality headset. Emotional valence during the task was assessed using both self-report questionnaires and a custom threshold-based algorithm that classified responses as positive or negative. Spectral features were extracted from the EEG recordings and subsequently used to train a k-nearest neighbors (KNN) classifier. The model was trained on data from 59 participants and evaluated on an independent test set of 25 participants, achieving an accuracy of 81.2% for positive valence (valence = 1) and 66.7% for negative valence (valence = 0). Beyond empirical analysis, we also explored mathematical models capable of reproducing signals with spectral characteristics comparable to those observed in the experimental data, thereby providing a framework for linking theoretical modeling with empirical findings. Two important remarks are necessary when interpreting these results. First, the VR-based scenarios, although effective in creating controlled experimental conditions, cannot fully replicate the affective dynamics of real-life situations. For example, the proportion of participants reporting anger was only 3.4%, which is considerably lower than would be expected in naturalistic contexts or even when viewing emotionally evocative film material [33].. This discrepancy likely reflects the original purpose of the VR scenarios, which were designed for employee training rather than affect induction, leading participants to adopt a more analytical than experiential stance during the task. Second, the EEG recordings were limited to four electrodes, which restricts the spatial resolution of the neural data. Despite these constraints, the classification algorithm performed robustly, achieving a good accuracy in distinguishing emotional valence, underscoring both the feasibility and the potential of combining EEG-based spectral features with machine learning for dynamic emotion recognition, see [34]. Our objective is to construct simplified mathematical models capable of reproducing signals indicative of neural electrical activity and corresponding emotional states. Future work will involve a systematic comparison between these simulated signals and empirical EEG recordings, as well as the integration and analysis of data acquired from EmotiBit devices.
@@
This study was part of the emergent projet MME funded by the Normandie Région. We thank also the Crédit Agricole Paris and the Hudson School of Mathematics for material support.

### v3

@@ 4 Discussion
In this study, we analyzed EEG data from a cohort of 87 participants exposed to a stressful work-related scenario presented through a virtual reality headset. Emotional valence during the task was assessed using both self-report questionnaires and a custom threshold-based algorithm that classified responses as positive or negative. Spectral features were extracted from the EEG recordings and subsequently used to train a k-nearest neighbors (KNN) classifier. The model was trained on data from 59 participants and evaluated on an independent test set of 25 participants, achieving an accuracy of 81.2% for positive valence (valence = 1) and 66.7% for negative valence (valence = 0). Beyond empirical analysis, we also explored mathematical models capable of reproducing signals with spectral characteristics comparable to those observed in the experimental data, thereby providing a framework for linking theoretical modeling with empirical findings. Two important remarks are necessary when interpreting these results. First, the VR-based scenarios, although effective in creating controlled experimental conditions, cannot fully replicate the affective dynamics of real-life situations. For example, the proportion of participants reporting anger was only 3.4%, which is considerably lower than would be expected in naturalistic contexts or even when viewing emotionally evocative film material [57].. This discrepancy likely reflects the original purpose of the VR scenarios, which were designed for employee training rather than affect induction, leading participants to adopt a more analytical than experiential stance during the task. Second, the EEG recordings were limited to four electrodes, which restricts the spatial resolution of the neural data. Despite these constraints, the classification algorithm performed robustly, achieving a good accuracy in distinguishing emotional valence, underscoring both the feasibility and the potential of combining EEG-based spectral features with machine learning for dynamic emotion recognition, see [58]. Then we discussed simplified mathematical models capable of reproducing signals indicative of neural electrical activity and corresponding emotional states. Future work will involve a systematic comparison between these simulated signals and empirical EEG recordings, as well as the integration and analysis of data acquired from EmotiBit devices.
@@
This study was part of the emergent projet MME funded by the Normandie Région. We thank also the Crédit Agricole Ile de France and the Hudson School of Mathematics for material support.